OpenStudy (anonymous):

if a man wants to calculate the WEIGHT of a perfect ball of metal, he would use Volume=(4π*R*R*R)/3 then multiply the answer times the density of a given material. The question is, how would you write the formula to reverse this process. All you know is the overall weight of the ball. Tell me how many inches a 10 pound ball is. (Density of metal is 0.409lbs per cube inch)

OpenStudy (cwrw238):

w = (4/3)* pi* r^3 * d

OpenStudy (anonymous):

WOW that was fast, than you@ Its been driving me nuts lol

OpenStudy (cwrw238):

what do u mean by how many inches? do you mean the radius r?

OpenStudy (anonymous):

Yes The problem is in there too, 0.409 lbs per square inch, and its a 10 pound ball... so I got to structure this to where I can tell what radius or diameter the ball is.

OpenStudy (anonymous):

excuse me, 0.409 lbs per CUBIC inch lol

OpenStudy (cwrw238):

rearranging the formula r^3 = w / ( 4/3 * pi * d)

OpenStudy (klimenkov):

\[\text{weight}=\text{volume}\times \text{density}\]\[\text{volume}=\frac {\text{weight}} {\text{density}}\]\[\text{volume}=\frac {4} {3} \pi R^3\]

OpenStudy (anonymous):

Me thinks too that the metal is lead, the specific gravity is about that of lead from what I can tell now, its bonus points lol I got that figured out,

OpenStudy (cwrw238):

r^3 = 10 / (4/3 * 3.14 * 0.409) r is cube root of this

OpenStudy (anonymous):

ok thank you, that looks right what I needed there